OpenStudy (anonymous):

MEDAL!!!!! The temperature of a chemical reaction ranges between 20 degrees Celsius and 160 degrees Celsius. The temperature is at its lowest point when t = 0, and the reaction completes 1 cycle during an 8-hour period. What is a cosine function that models this reaction?

OpenStudy (anonymous):

answer choices f(t) = -90 cos pi/4t +70 f(t) = -70 cos pi/4t +90 f(t) = 70 cos 8t +90 f(t)= 90 cos 8t +70

OpenStudy (anonymous):

Let's figure out the amplitude of the cosine function first. Given that the min is 20 degrees and max is 160 degrees, what is the amplitude?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

Amplitude should be calculated as \[A = \frac{ Max - Min }{ 2 }\]

OpenStudy (anonymous):

70

OpenStudy (anonymous):

Correct, so A and D are gone.

OpenStudy (anonymous):

answer is c

OpenStudy (anonymous):

So now calculate omega such that \[70 \cos(\omega t) + shift\]

OpenStudy (anonymous):

We are given the period 8.

OpenStudy (anonymous):

\[\omega = \frac{ 2 \pi}{ k }\]

OpenStudy (anonymous):

Where k = 8 (period)

OpenStudy (anonymous):

What is omega in this case?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

\[\omega = \frac{ 2\pi }{ 8 } = \frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

answer is b

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

can i ask another question

OpenStudy (anonymous):

Compare the functions below: f(x) = −3 sin(x − π) + 2 g(x) x y 0 8 1 3 2 0 3 −1 4 0 5 3 6 8 h(x) = (x + 7)^2 − 1

OpenStudy (anonymous):

which function has the smallest minimum

OpenStudy (anonymous):

Let's look at the f(x). We can quickly determine the minimum of the function by looking at its amplitude.

OpenStudy (anonymous):

Without the shift, it would be -3 right? Combining the vertical shift of 2, the minimum is -3+2=-1.

OpenStudy (anonymous):

i believe that h(x) has the smallest minimum value

OpenStudy (anonymous):

The minimum of g(x) is clearly shown in the table as -1.

OpenStudy (anonymous):

h(x) minimum is also -1

OpenStudy (anonymous):

Now, h(x) is a parabola. A parabola is normally centered at the origin, but it has been shifted downwards by -1.

OpenStudy (anonymous):

Correct, all of them have the same minimum.

OpenStudy (anonymous):

thanks